Two blocks of mass m = 10.7 kg each are fastened to the ceiling of an elevator. The elevator accelerates upward at a = 1.58 m/s2. Find the tension in the bottom rope. Find the tension in the top rope.
Answer #1

Solution :

Given :

m = 10.7 kg

a = 1.58 m/s2

Tension in the Bottom rope = m ( g + a ) = (10.7 kg)( 9.8 m/s2 +1.58 m/s2) = 121.766 N

Tension in the Top rope = 2 m ( g + a ) = 2(10.7 kg)(9.8 m/s2 + 1.58m/s2) = 243.532 N